A gas mixture containing oxygen, nitrogen, and carbon dioxide has a pressure of 250 mmHg. if the partial pressure of oxygen is 50 mmHg and the partial pressure of nitrogen is 175 mmHg, what is the partial pressure of the carbon dioxide?


1
Expert's answer
2022-02-04T01:30:03-0500

P total = P oxygen + P nitrogen + P CO2

250 = 50 + 175 + P CO2

P CO2 = 250-225 = 25

Partial pressure of CO2 ;

= 25mmHg